OpenStudy (anonymous):

Which of the following is a solution to the equation? 35x - 2(x - 30) = 48 + 3x Answer options: 40 -4 -.4 4 PLEASE PLEASE PLEASE help!

OpenStudy (anonymous):

-.4

OpenStudy (anonymous):

can you explain it @begzat :) I just want to double check :)

OpenStudy (anonymous):

35x-2x+60=48+3x 30x+12=0 x=-12/30=-0.4

OpenStudy (anonymous):

thanks! @isaacme

OpenStudy (anonymous):

ur wellcome